OpenStudy (anonymous):

Part 1: Find the perimeter of a rectangular object which has a length √128 of feet and a width of √200 feet. (3 points) Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (ghazi):

part 1 ..perimeter of rectangle = 2(length+ breadth) just put the values

OpenStudy (ghazi):

perimeter = total length along the sides = L+W+L+W =2(L+W)

OpenStudy (ghazi):

\[2 (\sqrt {128}+\sqrt {200})\]

OpenStudy (ghazi):

object around you that is rectangular , like study table, A4 sheet etc

OpenStudy (anonymous):

im still kind of lost... :/

OpenStudy (ghazi):

\[2(\sqrt{128}+\sqrt{200})= 2(8 \sqrt 2+ 10 \sqrt 2)= 2(18 \sqrt 2)=36 \sqrt 2\]

OpenStudy (ghazi):

@TiaW clear?

OpenStudy (anonymous):

Part 2: you basically add all the sides of a shape to get its perimeter. so for rectangle you both lengths and both widths to get its perimeter. this gives the equation 2(length + breadth) this is the shorter way of finding out the perimeter of a rectangle the longer way is l+ l +w + w =P of rectangle and then you just substitute the value in whatever equation you prefer. and this is how by calculating you reach the answer, the perimeter of the rectangle.

OpenStudy (anonymous):

Oook thanks @ghazi & @Bhagyashree

OpenStudy (anonymous):

@ghazi in the first on u put breadth?? whats that

OpenStudy (ghazi):

breadth= width :)

OpenStudy (anonymous):

it is the same as width

OpenStudy (anonymous):

Ooo thnxx

OpenStudy (ghazi):

:)

OpenStudy (anonymous):

Why does it not times the 2 (the radicand) as well?

OpenStudy (magbak):

What object do you think it is

OpenStudy (ghazi):

it's not an object i guess, because it's something two dimensional which is hard to analyse in the form of something 3D